Decide which of the two methods for graphing lines------the intercept method or the splope intercept method-----is more appropricate for graphing the equation
4x-9=39

Found 2 solutions by tutorcecilia, checkley71:
Answer by tutorcecilia(2152)

4x-9=39
4x-9+9=39+9
4x=48
4x/4=48/4
x=12
.
The intercept method is more appropriate because there is no y-intercept for the slop-intercept method.

Answer by checkley71(8403)

4X-9=39
4X=39+9
4X=48
X=48/4
X=12 ANSWER.
FOR THIS EXAMPLE THE INTERCEPT IS BETTER AS THIS LINE IS A VERTICLE LINE THROUGH THE POINT (12,0)